Assume that the vector field A= 5xyax + xy2ay + 4zaz is defined in a region 1≤ x ≤ 3, −2 ≤ y ≤ 4 and −1≤ z ≤ 2. Find the value of the closed-surface integral ∫∫ A.ds over the surface of this region.
1
Expert's answer
2020-03-17T10:02:38-0400

The vector field is


"\textbf{A}= 5xy\textbf{x} + xy^2\textbf{y} + 4z\textbf{z}."

The value of the closed-surface integral "\int\int AdS" over the surface of this region can be found with Gauss's theorem:


"\Phi=\iint\textbf{A}\text{ d}S=\iiint\text{ div}\textbf{A}\text{ d}V,\\
\space\\
\text{ div}\textbf{A}=\frac{\partial A_x}{\partial x}+\frac{\partial A_y}{\partial y}+\frac{\partial A_z}{\partial z}=\\
\space\\
=\frac{\partial 5xy}{\partial x}+\frac{\partial xy^2}{\partial y}+\frac{\partial 4z}{\partial z}=\\
\space\\
=5y+2xy+4.\\
\space\\
\iiint\text{ div}\textbf{A}\text{ d}V=\iiint(5y+2xy+4)\text{d}x\text{d}y\text{d}z=\\
\space\\
=\int_{-1}^2\Bigg[\int_{-2}^4\Bigg(\int^3_1(5y+2xy+4)\text{d}x\Bigg)\text{d}y\Bigg]\text{d}z=\\
\space\\
=\int_{-1}^2\Bigg[\int_{-2}^4\Bigg((5xy+x^2y+4x)\bigg|^3_1\Bigg)\text{d}y\Bigg]\text{d}z=\\
\space\\
=\int_{-1}^2\Bigg[\int_{-2}^4(18y+8)\text{d}y\Bigg]\text{d}z=\\
\space\\
=\int_{-1}^2\Bigg[(y^2+8y)\bigg|^4_{-2}\Bigg]\text{d}z=\\
\space\\
=\int_{-1}^260\text{ d}z=60z\bigg|^2_{-1}=180."